### v2

# On Picard’s Theorem Via Nevanlinna Theory

Abstract.
We study a long-standing Picard’s problem for non-compact complete Kähler manifolds with non-negative Ricci curvature. When the manifold carries a positive Green function, a positive answer is given to the Picard’s problem, i.e., it is showed that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values.
Key words and phrases:
Nevanlinna theory; defect relation; Picard type theorem2010 Mathematics Subject Classification:
32H30; 32H25Contents
1. Introduction
1.1. Motivations and Techniques
It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values. All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold. Due to S. T. Yau [44], each holomorphic function on a non-compact complete Kähler manifold (Hermitian manifold, more general) with non-negative Ricci curvature has Liouville property. So, a natural problem is that does Picard’s little theorem still hold on such a manifold? It is called the Picard’s problem. To our knowledge, this is a long-standing problem. In what follows, we shall give an accurate statement.
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature, see examples for such a manifold in Sha-Yang [34] and Tian-Yau [39, 40]. The well-known Picard’s problem asks that
Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
The first result in this direction could be traced back to 1970. S. Kobayashi [20] obtained some Picard-type theorems for holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity. As a special consequence, he showed that
Theorem A (Picard’s Theorem, Kobayashi).
Every meromorphic function on on which, a complex Lie group acts transitively, must be a constant if it omits distinct values.
Kobayashi’s result is limited to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [17] presented some Picard-type theorems on harmonic mappings of bounded dilatation on a locally flat Riemannian manifold, refer to N. Petridis [27] also. In particular, they showed that
Theorem B (Picard’s Theorem, Goldberg-Ishihara-Petridis).
Assume that is locally flat. Every holomorphic mapping of bounded dilatation must be a constant mapping.
It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation. The first result without any restrictions for is due to S. T. Yau [44], who considered the Liouville’s problem based on a method of gradient estimates at the same time, which is viewed as a weaker version of the Picard’s problem. He showed that
Theorem C (Liouville’s Theorem, Yau).
Every holomorphic function on must be a constant if it is bounded.
By S. T. Yau [44], Liouville’s theorem still holds for all harmonic functions on (see Cheng-Yau [9] also). The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem. Until 2010, A. Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate. In the following, we introduce his basic ideas. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with Fubini-Study metric Let be a meromorphic function. Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
He showed that
Theorem D (Picard’s Theorem, Atsuji).
Let be a meromorphic function. Assume that satisfies the growth condition
where is a geodesic ball centered at a fixed point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s trick [2] (see [11] also) is employing the probabilistic approach (via Brownian motions), who introduced the so-called Nevanlinna-type functions and in which is the time of a Brownian motion on By using Itô’s formula (see, e.g., [18]) and estimates of curvatures, he established a heat diffusion version of Second Main Theorem, which leads to the above Theorem D. It is apparent that is stochastically complete (or is conservative) since has non-negative Ricci curvature by Grigor’yan’s criterion [15]. However, in order to make and meaningful in spirit of Nevanlinna’s settings, the following conditions are necessary:
for ;
as ;
if omits
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition is so strong that few meromorphic functions can satisfy it.
In the present paper, we wish to settle the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature. As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R. Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925. During this period, rich results have been achieved, we may refer the reader to L. V. Ahlfors [4], H. Cartan [6], Carlson-Griffths-King [7, 16], J. Noguchi [23, 24], E. I. Nochka [25], M. Ru [28, 29], B. Shiffman [30], B. Shabat [31], F. Sakai [32, 33], W. Stoll [37, 38], P. Vojta [36], H. Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and their references therein. To our knowledge, almost all theorems from Nevanlinna theory are based only on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we know very little if the domain manifolds are not non-positively curved, since we know little about how to estimate Green functions (satisfying Dirichelet boundary condition) for bounded domains under a Ricci curvature condition.
In 2023, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature. As a consequence, the author gave a positive answer to the Picard’s problem if is of maximal volume growth saying that
where denotes the Riemannian volume of geodesic ball centered at a fixed reference point with radius in
The original technique in [14] is to construct a family of relatively compact domains exhausting through the global Green functions for based on an asymptotic estimate of minimal positive global Green function obtained by Colding-Minicozzi [8]. With an optimal estimate for local Green function for in terms of integral forms, the author (see [14]) established a Second Main Theorem of meromorphic mappings on Let us introduce the main contributions or key techniques in [14]. Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) obtained the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of the asymptotic estimate of the author defined the domain:
Thus, the Green function of for with a pole at satisfying Dirichlet boundary condition can be written as
According to [14], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma can be established. The situation where is trivial for that is conformally equivalent to under two assumptions of non-negative Ricci curvature and maximal volume growth (see [21]). By employing the standard arguments, the author established a Second Main Theorem of meromorphic mappings from into a complex projective manifold, which gives a Picard’s theorem:
Theorem E (Picard’s Theorem, Dong).
Assume that is of maximal volume growth. Every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [14] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no asymptotic estimates for if is not of maximal volume growth.
In this paper, we will give a positive answer under a much weaker volume growth condition:
which is equivalent to that carries a positive global Green function. Refer to N. Varopoulos [41, 42] (or see Li-Tam-Wang [19, 21]), there exists uniquely a minimal positive global Green function of for denoted by Our original method is the construction of domain and the application of Li-Yau’s estimates on heat kernels and Green functions.
According to the estimate for obtained by Li-Yau [22], there exist constants such that
Re-define by
In further, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and (see Lemma 4.1): for we have
which plays an important role in the establishment of Nevanlinna theory in this paper.
In the study of Nevanlinna theory, Negative Curvature Method and Logarithmic Derivative Lemma Method are considered as two standard research approaches. In our investigations, Logarithmic Derivative Lemma Method is used. In doing so, it is necessary to establish the so-called Calculus Lemma, which turns out to be a workable road to the Logarithmic Derivative Lemma.
Let us introduce the main techniques: Gradient Estimation (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimation. For any there exists such that
holds for all
The above gradient estimate of gives an estimate of the harmonic measure on with respect to : for any there exists such that (see Corollary 4.3)
holds for all
Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Let be a nonconstant meromorphic function on The Calculus Lemma can lead to Logarithmic Derivative Lemma (see Theorem 5.6): for any there exists a subset of finite Lebesgue measure such that
holds for all outside
1.2. Main results
Let be a complex projective manifold with over which we can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper. Assume that carries a positive global Green function, i.e., it satisfies the volume growth condition:
Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined as
The first main result is the following Second Main Theorem.
Theorem 1.1 (=Theorem 6.1).
Let be a differentiably non-degenerate meromorphic mapping. Let be a reduced divisor of simple normal crossing type. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
For a divisor the simple defect of with respect to is defined by
Again, put
Theorem 1.1 gives a defect relation:
Corollary 1.2 (=Corollary 6.4).
Assume the same conditions as in Theorem Then
Let be the scalar curvature of Set
We obtain a Picard’s theorem:
Corollary 1.3 (=Corollary 6.11).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values. Actually, this theorem gives a quantitative solution to the Picard’s problem.
2. Some Facts from Differential Geometry and Geometric Analysis
2.1. Volume Comparison Theorem
A space form is defined as a complete Riemannian manifold with constant sectional curvature. Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Denote by the Riemannian volume of a geodesic ball centered at with radius
Bishop-Gromov (see, e.g., [5]) gave an upper bound of :
Theorem 2.1 (Volume Comparison Theorem).
If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
holds for all
In particular, if then we obtain:
Corollary 2.2.
Assume that is non-compact. If then we have
holds for all where is the volume of a unit ball in
Calabi-Yau (see, e.g., [35]) gave a lower bound of :
Theorem 2.3.
Assume that is non-compact. If then has an infinite volume. More precisely, there exists a constant such that
holds for all
2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of for is the minimal positive fundamental solution of the following heat equation
Fix any reference point Let be the Riemannian distance function of from
Li-Yau [22] gave an estimate for :
Theorem 2.4.
Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
holds for all
Set
If the right-hand integral converges , then defines a minimal positive global Green function of for with a pole at i.e.,
where is the Dirac’s delta function with a pole at It is evident that such a Green function is also unique if (1) is satisfied.
In further, Li-Yau [22] obtained an estimate for :
Theorem 2.5.
Assume that has non-negative Ricci curvature. If exists, then there exist constants depending only on such that
holds for all
3. Nevanlinna’s Functions and First Main Theorem
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature of complex dimension whose Kähler form is defined by
in a local holomorphic coordinate Fix a reference point Let stand for the geodesic ball centered at with radius in Denote by the Riemannian volume of
3.1. Construction of
Assume that is satisfied the condition of volume growth:
It implies that is non-parabolic and there exists a unique minimal positive global Green function for satisfying
where denotes the Laplace-Beltrami operator and is the Dirac function with a pole at Let be the Riemannian distance function of from According to Li-Yau [22], there exist constants such that
Define
It is evident that is relatively compact for any and the sequence exhausts if According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
which is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Moreover, we denote by the harmonic measure on with respect to which is defined by
where is the inward normal derivative on is the Riemannian area element of
3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form where
Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
where is the volume element of
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential By the Kählerity of we have
Locally, write where is a local holomorphic frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [7]), it leads to
in the sense of currents. So, we obtain the alternative expressions of and as follows
and
3.3. First Main Theorem
To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) which states that
Lemma 3.1 (Jensen-Dynkin formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Jensen-Dynkin formula to then we are led to
Therefore, we are led to that
Theorem 3.2 (First Main Theorem).
Assume that Then
4. Gradient Estimates of Green Functions
Let be a non-compact complete Kähler manifold of complex dimension satisfying
Theorem 4.1.
We have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since
then we obtain
Let denote the gradient operator on We obtain an estimate of upper bounds of as follows:
Theorem 4.2.
For any there exists such that
holds for all where is given by
Proof.
It yields from (2) that
On the other hand, we have
Whence, it concludes that
where is the cut locus of and is the inward normal derivative on By
we see that for any there exists such that
holds for all Since has measure 0, with the aid of the continuity of on and on one then has the theorem proved. ∎
As a result, we obtain an estimate for upper bounds of harmonic measure on with respect to as follows:
Corollary 4.3.
For any there exists such that
holds for all where is given by
Proof.
5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-compact complete Kähler manifold of complex dimension satisfying
5.1. Calculus Lemma
We need the following Borel’s lemma (see, e.g., [24]):
Lemma 5.1 (Borel Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
holds for all outside
Proof.
The conclusion is true clearly for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
exists, here is allowed. If then Set
Note that exists for almost every Then, we have
This completes the proof. ∎
Now, we give the Calculus Lemma:
Theorem 5.2 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Invoking Theorem 4.1, we obtain
Set
A simple computation leads to
In further, we have
Now, we apply Borel’s lemma to the left hand side of the above equality for twice: one is to and the other is to we conclude that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
holds for all where is given by (2). Combining the above, we have
holds for all outside Therefore, we have the theorem proved by setting
As a result, we conclude that
Corollary 5.3.
Let be a locally integrable function on Assume that is locally bounded at Assume that is of maximal volume growth. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on The norm of the gradient is defined by
in a local holomorphic coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
Then, it leads to
Lemma 5.4.
We have
Proof.
Note that
Whence, it concludes from Fubini’s theorem that
Lemma 5.5.
Assume that Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
The concavity of implies that
By this with Corollary 5.3, we show that for any there exists a subset of finite Lebesgue measure such that
holds for all outside Disturbing such that is absorbed. The proof is completed. ∎
Define
Next, we establish the following Logarithmic Derivative Lemma:
Theorem 5.6 (Logarithmic Derivative Lemma).
Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Note that
Using Lemma 5.5 and disturbing we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside ∎
6. Second Main Theorem and Defect Relation
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature satisfying
Set
6.1. Second Main Theorem
Let be a reduced divisor of simple normal crossing type. We write which is an irreducible decomposition of Equip every holomorphic line bundle with a Hermitian metric which induces a Hermitian metric on Since then we can assume that As a result, gives a volume form on Pick such that and Moreover, we define a singular volume form
on Set
It is clear that
In further, we have
in the sense of currents, where is the reduced divisor of and is the Ricci form of Therefore, it yields that
Let us establish a Second Main Theorem as follows:
Theorem 6.1 (Second Main Theorem).
Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
The argument of proof is standard (see [14]), but for the completeness of this paper, we still give a proof. Since has only simple normal crossings, then there exist a finite open covering of and finitely many rational functions on such that are holomorphic on satisfying that
In addition, we can require that for On write
where is a positive smooth function on Let be a partition of the unity subordinate to Set
Again, put On we have
Fix a one can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
Put
Then, we have and at :
Define a non-negative function on by
Again, set for Then, at :
That is, at :
Combining the above, we are led to
on Note that is bounded on whence it yields from that
By Jensen-Dynkin formula
Combining (6) with (7) and using Theorem 5.6, we obtain
Using Theorem 5.2 and (5), for any there exists a subset of finite Lebesgue measure such that
holds for all outside Thus, we conclude that
for all outside Combining this with (6.1), we prove the theorem. ∎
Let be a compact Riemann surface of genus Let be the holomorphic line bundle over defined by a point Note that is independent of the choice of It yields from that
Corollary 6.2.
Let be distinct points in Let be a nonconstant meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the hyperplane line bundle over and whose Chern form is given by It yields from that
Corollary 6.3.
Let be hyperplanes in general position in Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
6.2. Defect Relation
For any divisor the defect and simple defect of with respect to are defined respectively by
Using the First Main Theorem, we see that
For two holomorphic line bundles over we define (see [7])
Corollary 6.4 (Defect Relation).
Assume the same conditions as in Theorem Then
Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Denote by the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
where
Using the definition of scalar curvature, we have
In further, it yields that
Corollary 6.5.
Let be a meromorphic mapping. If there exists a small number such that satisfies the growth-curvature condition
for sufficiently large then must be differentiably degenerate.
Proof.
Assume on the contrary that is differentiably non-degenerate. Then, we have as The growth-curvature condition leads to
Thus, we obtain due to Corollary 6.4, but which contradicts with ∎
Treat a compact Riemann surface of genus we have:
Corollary 6.6.
Let be a nonconstant meromorphic mapping. Let be distinct points in Then
For a meromorphic mapping
we have
with respect to metrics
Corollary 6.7.
Let be a differentiably non-degenerate meromorphic mapping. Let be hyperplanes in general position in Then
A meromorphic function on can be seen as a meromorphic mapping Set
We obtain a Picard’s theorem:
Corollary 6.8 (Picard Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
6.3. Picard’s Theorem
In the following, we investigate a more exact form of the Picard’s theorem given in Corollary 6.8. On we introduce the spherical distance
Let be a meromorphic function on which is regarded as a meromorphic mapping
In terms of the spherical distance, the proximity function of with respect to can be defined as
We use the Ahlfors-Shimizu’s characteristic function of :
Moreover, we define
It is not hard to show the First Main Theorem:
Theorem 6.9 (Second Main Theorem).
Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
By
in the sense of distributions, we deduce that
Combining this with Corollary 6.3, it gives immediately
We re-define the simple defect of with respect to by
Set
Corollary 6.10 (Defect Relation).
Assume the same conditions as in Theorem Then
Consequently, we derive a more exact Picard’s theorem:
Corollary 6.11 (Picard’s Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than

### v5

@@
 Abstract.
-We study a long-standing Picard’s problem for non-compact complete Kähler manifolds with non-negative Ricci curvature.
-When the manifold carries a positive Green function, a positive answer is given to the Picard’s problem, i.e., it is showed that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values.
+We consider the Picard’s problem via Nevanlinna theory for complete Kähler manifolds with non-negative Ricci curvature.
+When the manifold is non-parabolic, which means that it carries a positive global Green function, we give a positive answer to the Picard’s problem.
+That is, we confirm that every meromorphic function on such a (non-compact) manifold reduces to a constant if it omits three distinct values.
 Key words and phrases:
@@
 T.
-Yau [44], each holomorphic function on a non-compact complete Kähler manifold (Hermitian manifold, more general) with non-negative Ricci curvature has Liouville property.
+Yau [44], each holomorphic function on a non-compact complete Kähler manifold (or Hermitian manifolds in general) with non-negative Ricci curvature has Liouville property.
 So, a natural problem is that does Picard’s little theorem still hold on such a manifold?
-It is called the Picard’s problem.
-To our knowledge, this is a long-standing problem.
-In what follows, we shall give an accurate statement.
+We call it the Picard’s problem.
+To our knowledge, it is still an unsolved problem.
+In what follows, we would like to give a precise statement.
 Let be a non-compact complete Kähler manifold with non-negative Ricci curvature, see examples for such a manifold in Sha-Yang [34] and Tian-Yau [39, 40]. The well-known Picard’s problem asks that
 Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
-The first result in this direction could be traced back to 1970.
-S.
-Kobayashi [20] obtained some Picard-type theorems for holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity.
+Remark.
+Let be the universal covering, where is the unit disc in It is not very hard to confirm that the Picard’s problem can be given a positive answer using Yau’s theorem (see Theorem C below), if any holomorphic mapping lifts to a holomorphic mapping via i.e., Using the criterion for the existence of the liftings, can be lifted to if and only if for some such that Since is simply-connected, then is trivial.
+This implies that such a lifting exists if and only if is trivial.
+However, is often not a trivial group, so one cannot lift to in general.
+The first result in this direction could be traced back to 1970, S.
+Kobayashi [20] proved some Picard-type theorem for a holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity.
 As a special consequence, he showed that
@@
 T.
-Yau [44], who considered the Liouville’s problem based on a method of gradient estimates at the same time, which is viewed as a weaker version of the Picard’s problem.
+Yau [44], who considered the Liouville’s problem based on the gradient estimation method at the same time, which is viewed as a weaker version of the Picard’s problem.
 He showed that
@@
 T.
-Yau [44], Liouville’s theorem still holds for all harmonic functions on (see Cheng-Yau [9] also).
+Yau [44], Liouville’s theorem works for all the harmonic functions on (see Cheng-Yau [9] also).
 The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem.
@@
 Unfortunately, the growth condition is so strong that few meromorphic functions can satisfy it.
-In the present paper, we wish to settle the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature.
+In the present paper, we shall consider the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature.
 As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R.
@@
 Vojta [36], H.
-Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and their references therein.
-To our knowledge, almost all theorems from Nevanlinna theory are based only on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we know very little if the domain manifolds are not non-positively curved, since we know little about how to estimate Green functions (satisfying Dirichelet boundary condition) for bounded domains under a Ricci curvature condition.
-In 2023, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature.
-As a consequence, the author gave a positive answer to the Picard’s problem if is of maximal volume growth saying that
+Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and references therein.
+To our knowledge, almost all the results in Nevanlinna theory are built just on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, because it is hard to estimate the local Green functions for bounded domains under other curvature conditions.
+Recently, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature by introducing a new method: global Green function method.
+Based on this method, a positive answer to the Picard’s problem can be given provided that is of maximal volume growth, i.e.,
 where denotes the Riemannian volume of geodesic ball centered at a fixed reference point with radius in
-The original technique in [14] is to construct a family of relatively compact domains exhausting through the global Green functions for based on an asymptotic estimate of minimal positive global Green function obtained by Colding-Minicozzi [8].
-With an optimal estimate for local Green function for in terms of integral forms, the author (see [14]) established a Second Main Theorem of meromorphic mappings on Let us introduce the main contributions or key techniques in [14].
-Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) obtained the asymptotic behavior of : there exists a constant such that
-where denotes the Riemannian distance function of from By means of the asymptotic estimate of the author defined the domain:
+The original technique in [14] is to construct a family of relatively compact domains exhausting by using a minimal positive global Green function for Based on the asymptotic estimate of obtained by Colding-Minicozzi [8], we can give an optimal estimate for the local Green function for in terms of integral forms.
+As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce the main contribution and key technique in [14].
+Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) described the asymptotic behavior of : there exists a constant such that
+where denotes the Riemannian distance function of from By means of this asymptotic estimate, the author defined a family of relatively compact domains:
 Thus, the Green function of for with a pole at satisfying Dirichlet boundary condition can be written as
-According to [14], for any there exists such that
-holds for all satisfying Based on this estimate, the Calculus Lemma can be established.
-The situation where is trivial for that is conformally equivalent to under two assumptions of non-negative Ricci curvature and maximal volume growth (see [21]).
-By employing the standard arguments, the author established a Second Main Theorem of meromorphic mappings from into a complex projective manifold, which gives a Picard’s theorem:
+According to Lemma 3.5 in [14], for any there exists such that
+holds for all satisfying Based on this estimate, the Calculus Lemma can be derived.
+When is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [21]).
+So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [14], Theorem 3.10), which gives a Picard’s theorem:
 Theorem E (Picard’s Theorem, Dong).
@@
 However, the technique in [14] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no asymptotic estimates for if is not of maximal volume growth.
-In this paper, we will give a positive answer under a much weaker volume growth condition:
+In this paper, we will consider a weaker volume growth condition:
 which is equivalent to that carries a positive global Green function.
@@
 Varopoulos [41, 42] (or see Li-Tam-Wang [19, 21]), there exists uniquely a minimal positive global Green function of for denoted by Our original method is the construction of domain and the application of Li-Yau’s estimates on heat kernels and Green functions.
-According to the estimate for obtained by Li-Yau [22], there exist constants such that
+By Li-Yau’s estimate [22], there exist constants such that
 Re-define by
@@
 1.2. Main results
-Let be a complex projective manifold with over which we can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper.
-Assume that carries a positive global Green function, i.e., it satisfies the volume growth condition:
-Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined as
-The first main result is the following Second Main Theorem.
+Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper.
+Assume that is non-parabolic, by which we mean that carries a positive global Green function, i.e., it satisfies the volume growth condition
+Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
+The main result is the following Second Main Theorem.
 Theorem 1.1 (=Theorem 6.1).
-Let be a differentiably non-degenerate meromorphic mapping. Let be a reduced divisor of simple normal crossing type. Then for any there exists a subset of finite Lebesgue measure such that
+Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
+Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
+If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
@@
 Assume the same conditions as in Theorem Then
-Let be the scalar curvature of Set
+Let denote the scalar curvature of Set
 We obtain a Picard’s theorem:
@@
 Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
-Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values. Actually, this theorem gives a quantitative solution to the Picard’s problem.
+Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values.
 2. Some Facts from Differential Geometry and Geometric Analysis
@@
 A space form is defined as a complete Riemannian manifold with constant sectional curvature.
-Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Denote by the Riemannian volume of a geodesic ball centered at with radius
-Bishop-Gromov (see, e.g., [5]) gave an upper bound of :
+Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Let denote the Riemannian volume of a geodesic ball centered at with radius
+The classical volume comparison theorem by Bishop-Gromov (see [5]) can be stated as follows.
 Theorem 2.1 (Volume Comparison Theorem).
@@
 holds for all where is the volume of a unit ball in
-Calabi-Yau (see, e.g., [35]) gave a lower bound of :
+When Calabi-Yau (see [35]) obtained a lower bound of
 Theorem 2.3.
-Assume that is non-compact. If then has an infinite volume. More precisely, there exists a constant such that
+Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
 holds for all
 2.2. Estimates of Heat Kernels
-Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of for is the minimal positive fundamental solution of the following heat equation
+Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of is the minimal positive fundamental solution of the following heat equation
 Fix any reference point Let be the Riemannian distance function of from
-Li-Yau [22] gave an estimate for :
 Theorem 2.4.
@@
 where is the Dirac’s delta function with a pole at It is evident that such a Green function is also unique if (1) is satisfied.
-In further, Li-Yau [22] obtained an estimate for :
 Theorem 2.5.
@@
 3.1. Construction of
-Assume that is satisfied the condition of volume growth:
-It implies that is non-parabolic and there exists a unique minimal positive global Green function for satisfying
+Assume that is non-parabolic, which means that satisfies the volume growth condition
+Hence, there exists a unique minimal positive global Green function of for i.e.,
 where denotes the Laplace-Beltrami operator and is the Dirac function with a pole at Let be the Riemannian distance function of from According to Li-Yau [22], there exist constants such that
 Define
-It is evident that is relatively compact for any and the sequence exhausts if According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
+By as it is evident to see that is relatively compact for any and the sequence exhausts for According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
 which is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
@@
 Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
-where is the volume element of
-Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential By the Kählerity of we have
+where is the Riemannian volume element of
+Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential Since is Kählerian, then we have
 Locally, write where is a local holomorphic frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [7]), it leads to
@@
 3.3. First Main Theorem
-To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) which states that
+To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) as follows.
 Lemma 3.1 (Jensen-Dynkin formula).
@@
 4. Gradient Estimates of Green Functions
-Let be a non-compact complete Kähler manifold of complex dimension satisfying
+Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
 Theorem 4.1.
@@
 5. Calculus Lemma and Logarithmic Derivative Lemma
-Let be a non-compact complete Kähler manifold of complex dimension satisfying
+Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
 5.1. Calculus Lemma
@@
 In further, we have
-Now, we apply Borel’s lemma to the left hand side of the above equality for twice: one is to and the other is to we conclude that for any there exists a subset of finite Lebesgue measure such that
+Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
@@
 Lemma 5.5.
-Assume that Then for any there exists a subset of finite Lebesgue measure such that
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 Proof.
-The concavity of implies that
-By this with Corollary 5.3, we show that for any there exists a subset of finite Lebesgue measure such that
-holds for all outside Disturbing such that is absorbed. The proof is completed. ∎
 Define
@@
 Theorem 5.6 (Logarithmic Derivative Lemma).
-Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
@@
 6. Second Main Theorem and Defect Relation
-Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature satisfying
+Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
 Set
+where is the Ricci form of
+Let be a divisor, where are prime divisors. The reduced form of is defined as
 6.1. Second Main Theorem
@@
 We write which is an irreducible decomposition of Equip every holomorphic line bundle with a Hermitian metric which induces a Hermitian metric on Since then we can assume that As a result, gives a volume form on Pick such that and Moreover, we define a singular volume form
-on Set
+on Let
+be a differentiably non-degenerate meromorphic mapping. Set
 It is clear that
 In further, we have
-in the sense of currents, where is the reduced divisor of and is the Ricci form of Therefore, it yields that
-Let us establish a Second Main Theorem as follows:
+in the sense of currents, where is the reduced form of Thus, it yields that
+We establish a Second Main Theorem as follows.
 Theorem 6.1 (Second Main Theorem).
+Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
 Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
-Then for any there exists a subset of finite Lebesgue measure such that
+If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
@@
 Assume the same conditions as in Theorem Then
-Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Denote by the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
+Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Let denote the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
 where
@@
 We obtain a Picard’s theorem:
-Corollary 6.8 (Picard Theorem).
+Corollary 6.8 (Picard’s Theorem).
 Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
@@
 Theorem 6.9 (Second Main Theorem).
-Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
+Let be distinct values in Let be a nonconstant meromorphic function on If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside

### v6

@@
 Abstract.
-We consider the Picard’s problem via Nevanlinna theory for complete Kähler manifolds with non-negative Ricci curvature.
-When the manifold is non-parabolic, which means that it carries a positive global Green function, we give a positive answer to the Picard’s problem.
-That is, we confirm that every meromorphic function on such a (non-compact) manifold reduces to a constant if it omits three distinct values.
+We study the classical Picard’s problem for complete Kähler manifolds with non-negative Ricci curvature.
+Based on the global Green function method, we generalize the Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature.
+As a consequence, we give a positive answer to the Picard’s problem in the non-parabolic case, i.e., if the (non-compact) manifold is non-parabolic, then we show that every meromorphic function is a constant if it omits three distinct values.
 Key words and phrases:
-Nevanlinna theory; defect relation; Picard type theorem2010 Mathematics Subject Classification:
-32H30; 32H25Contents
+Nevanlinna theory; meromorphic mappings; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
+32H30; 32H25; 32A22Contents
 1. Introduction
 1.1. Motivations and Techniques
-It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values.
-All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold.
-Due to S.
+The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces will reduce to a constant, if it omits three distinct values.
+Historically, people have been committed to developing this theorem for a long time (see [4, 16, 18, 20, 21, 24, 31]).
+The study of Picard’s problem on manifolds may have originated from the celebrated work due to S.
 T.
-Yau [44], each holomorphic function on a non-compact complete Kähler manifold (or Hermitian manifolds in general) with non-negative Ricci curvature has Liouville property.
-So, a natural problem is that does Picard’s little theorem still hold on such a manifold?
-We call it the Picard’s problem.
-To our knowledge, it is still an unsolved problem.
-In what follows, we would like to give a precise statement.
-Let be a non-compact complete Kähler manifold with non-negative Ricci curvature, see examples for such a manifold in Sha-Yang [34] and Tian-Yau [39, 40]. The well-known Picard’s problem asks that
+Yau [48] on Liouville’s theorem, which proves the Liouville property of harmonic functions on non-negatively Ricci curved manifolds.
+Yau’s work has caused a classical problem: does Picard’s theorem hold on complete Kähler manifolds with non-negative Ricci curvature?
+It is well-known that Nevanlinna theory [30] is a great development of both Liouville’s theorem and Picard’s theorem.
+So, a further problem is, can Nevanlinna theory extend to complete Kähler manifolds with non-negative Ricci curvature?
+As far as we know, neither of them has been solved.
+Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see such examples in Sha-Yang [38] and Tian-Yau [43, 44]). The classical Picard’s problem states that
 Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
 Remark.
-Let be the universal covering, where is the unit disc in It is not very hard to confirm that the Picard’s problem can be given a positive answer using Yau’s theorem (see Theorem C below), if any holomorphic mapping lifts to a holomorphic mapping via i.e., Using the criterion for the existence of the liftings, can be lifted to if and only if for some such that Since is simply-connected, then is trivial.
-This implies that such a lifting exists if and only if is trivial.
-However, is often not a trivial group, so one cannot lift to in general.
-The first result in this direction could be traced back to 1970, S.
-Kobayashi [20] proved some Picard-type theorem for a holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity.
-As a special consequence, he showed that
-Theorem A (Picard’s Theorem, Kobayashi).
-Every meromorphic function on on which, a complex Lie group acts transitively, must be a constant if it omits distinct values.
-Kobayashi’s result is limited to certain complex manifolds which are acted on transitively by a complex Lie group.
-In 1975, Goldberg-Ishihara-Petridis [17] presented some Picard-type theorems on harmonic mappings of bounded dilatation on a locally flat Riemannian manifold, refer to N.
-Petridis [27] also.
-In particular, they showed that
-Theorem B (Picard’s Theorem, Goldberg-Ishihara-Petridis).
-Assume that is locally flat. Every holomorphic mapping of bounded dilatation must be a constant mapping.
-It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation.
-The first result without any restrictions for is due to S.
-T.
-Yau [44], who considered the Liouville’s problem based on the gradient estimation method at the same time, which is viewed as a weaker version of the Picard’s problem.
+Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem of Yau, we may have an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by By the mapping lifting theorem, admits a lifting if and only if
+for some such that Since is simply-connected, then the fundamental group is trivial.
+It implies that is necessarily a trivial group if such a lifting exists.
+However, is not trivial in most cases.
+Thus, we cannot lift to in general.
+An early result in this direction could be dated back to 1970, S. Kobayashi [24] obtained some Picard-type theorems for holomorphic mappings between complex manifolds. In particular, he showed that
+Theorem A (Kobayashi, 1970).
+If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
+Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group.
+In 1975, Goldberg-Ishihara-Petridis [21] established some Picard-type theorems on the locally flat manifolds (see N.
+Petridis [31] also).
+They showed that
+Theorem B (Goldberg-Ishihara-Petridis, 1975).
+If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant mapping.
+In 2010, A.
+Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by developing a Nevanlinna-type theory based on heat diffusion, Let us introduce his main result.
+Let be the Kähler form and Laplace-Beltrami operator of associated to respectively.
+One can endow with the Fubini-Study metric Let be a meromorphic function.
+The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
 He showed that
-Theorem C (Liouville’s Theorem, Yau).
-Every holomorphic function on must be a constant if it is bounded.
-By S.
-T.
-Yau [44], Liouville’s theorem works for all the harmonic functions on (see Cheng-Yau [9] also).
-The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem.
-Until 2010, A.
-Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate.
-In the following, we introduce his basic ideas.
-Let be the Kähler form and Laplace-Beltrami operator of associated to respectively.
-One can endow with Fubini-Study metric Let be a meromorphic function.
-Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
-He showed that
-Theorem D (Picard’s Theorem, Atsuji).
+Theorem C (Atsuji, 2010).
 Let be a meromorphic function. Assume that satisfies the growth condition
-where is a geodesic ball centered at a fixed point with radius in Then, must be a constant if it omits distinct values.
-Atsuji’s trick [2] (see [11] also) is employing the probabilistic approach (via Brownian motions), who introduced the so-called Nevanlinna-type functions and in which is the time of a Brownian motion on By using Itô’s formula (see, e.g., [18]) and estimates of curvatures, he established a heat diffusion version of Second Main Theorem, which leads to the above Theorem D.
-It is apparent that is stochastically complete (or is conservative) since has non-negative Ricci curvature by Grigor’yan’s criterion [15].
-However, in order to make and meaningful in spirit of Nevanlinna’s settings, the following conditions are necessary:
+for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
+Atsuji’s method [2] (see [12] also) is using stochastic calculus by Brownian motion, who introduced the so-called Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [22]), he established a heat diffusion version of the Second Main Theorem which leads to the above Theorem C.
+Note that is stochastically complete (i.e., is conservative), since has non-negative Ricci curvature due to Grigor’yan’s criterion [17].
+However, in order to make and meaningful in the sense of Nevanlinna theory, the following conditions are necessary:
 for ;
@@
 That is why a growth condition is needed in Atsuji’s theorem.
-Unfortunately, the growth condition is so strong that few meromorphic functions can satisfy it.
-In the present paper, we shall consider the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature.
-As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R.
-Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925.
-During this period, rich results have been achieved, we may refer the reader to L.
-V.
+Unfortunately, the growth condition seems so strong that there are only a few meromorphic functions satisfy it.
+The main purpose of the present paper is to generalize Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature.
+It is known that Nevanlinna theory in several complex variables (see [28, 33]) studies the value distribution of meromorphic mappings between complex spaces, which has been developed for a long time.
+During this period, fruitful results have been achieved.
+For example, the reader may refer to L.
 Ahlfors [5], H.
@@
 Noguchi [27, 28], E.
-I.
 Nochka [29], M.
@@
 Stoll [41, 42], P.
-Vojta [36], H.
-Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and references therein.
-To our knowledge, almost all the results in Nevanlinna theory are built just on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, because it is hard to estimate the local Green functions for bounded domains under other curvature conditions.
-Recently, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature by introducing a new method: global Green function method.
-Based on this method, a positive answer to the Picard’s problem can be given provided that is of maximal volume growth, i.e.,
-where denotes the Riemannian volume of geodesic ball centered at a fixed reference point with radius in
-The original technique in [14] is to construct a family of relatively compact domains exhausting by using a minimal positive global Green function for Based on the asymptotic estimate of obtained by Colding-Minicozzi [8], we can give an optimal estimate for the local Green function for in terms of integral forms.
-As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce the main contribution and key technique in [14].
-Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) described the asymptotic behavior of : there exists a constant such that
-where denotes the Riemannian distance function of from By means of this asymptotic estimate, the author defined a family of relatively compact domains:
-Thus, the Green function of for with a pole at satisfying Dirichlet boundary condition can be written as
+Vojta [40] and H.
+Wu [47], etc.; and also refer to [1, 2, 3, 12, 13, 14, 15].
+To our knowledge, almost all results in Nevanlinna theory are built on complex manifolds with a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, since it is difficult to estimate the local Green function for a general bounded domain, even for a geodesic ball, under other curvature conditions.
+Recently, the author [15] gave the first result on the Second Main Theorem in Nevanlinna theory on complete Kähler manifolds with non-negative Ricci curvature by introducing the global Green function method.
+Consequently, the author answered the Picard’s problem affirmatively provided that is of maximal volume growth:
+where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
+The original technique in [15] is to construct a family of relatively compact exhaustive domains of using the minimal positive global Green function for Based on an asymptotic estimate of obtained by Colding-Minicozzi [9], one can obtain two-sided bounds of the local Green function for in terms of integral forms.
+As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce his main result as follows.
+Colding-Minicozzi [9] (see [25] also) described the asymptotic behavior of : there exists a constant such that
+where denotes the Riemannian distance function of from By means of the asymptotic estimate, the author defined a family of relatively compact domains as follows
+for Hence, the Green function of for with a pole at satisfying Dirichlet boundary condition can be formulated as
 According to Lemma 3.5 in [15], for any there exists such that
-holds for all satisfying Based on this estimate, the Calculus Lemma can be derived.
-When is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [21]).
-So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [14], Theorem 3.10), which gives a Picard’s theorem:
-Theorem E (Picard’s Theorem, Dong).
-Assume that is of maximal volume growth. Every meromorphic function on must be a constant if it omits distinct values.
-However, the technique in [14] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no asymptotic estimates for if is not of maximal volume growth.
-In this paper, we will consider a weaker volume growth condition:
-which is equivalent to that carries a positive global Green function.
-Refer to N.
-Varopoulos [41, 42] (or see Li-Tam-Wang [19, 21]), there exists uniquely a minimal positive global Green function of for denoted by Our original method is the construction of domain and the application of Li-Yau’s estimates on heat kernels and Green functions.
-By Li-Yau’s estimate [22], there exist constants such that
-Re-define by
-In further, we have
-which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and (see Lemma 4.1): for we have
-which plays an important role in the establishment of Nevanlinna theory in this paper.
-In the study of Nevanlinna theory, Negative Curvature Method and Logarithmic Derivative Lemma Method are considered as two standard research approaches.
-In our investigations, Logarithmic Derivative Lemma Method is used.
-In doing so, it is necessary to establish the so-called Calculus Lemma, which turns out to be a workable road to the Logarithmic Derivative Lemma.
-Let us introduce the main techniques: Gradient Estimation (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
-Gradient Estimation. For any there exists such that
+holds for all satisfying Based on this estimate, the Calculus Lemma was established.
+For is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [25]).
+So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [15], Theorem 3.10), which gives an immediate result:
+Theorem D (Dong, 2023).
+If is of maximal volume growth, then every meromorphic function on must be a constant if it omits distinct values.
+However, the technique in [15] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no such asymptotic estimates for if is not of maximal volume growth.
+In this paper, we will generalize the Nevanlinna theory in the case that is a non-parabolic manifold, i.e., there exists a positive global Green function for Note that is non-parabolic if is of maximal volume growth.
+By N.
+Varopoulos [45, 46] (see Li-Tam-Wang [23, 25] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted also by Our original idea is the construction of domain and application of Li-Yau’s estimate on heat kernel and global Green function.
+Using Li-Yau’s estimate [26], there exist constants such that
+holds for all and For we redefine
+Then, we have
+which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and : for we have
+which plays an important role in the establishment of Nevanlinna theory in this paper (see Lemma 4.1).
+Negative Curvature Method and Logarithmic Derivative Lemma Method are very typical approaches in the study of Nevanlinna theory.
+In this paper, the Logarithmic Derivative Lemma Method is used.
+So, we need to establish the so-called Calculus Lemma which is a way to the Logarithmic Derivative Lemma.
+We will introduce two techniques in the paper: Gradient Estimates (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
+Gradient Estimates. For any there exists such that
 holds for all
-The above gradient estimate of gives an estimate of the harmonic measure on with respect to : for any there exists such that (see Corollary 4.3)
-holds for all
+The above gradient estimate yields an estimate of the harmonic measure on with respect to : for any there exists such that
+holds for all (see Corollary 4.3).
 Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
 holds for all outside
-Let be a nonconstant meromorphic function on The Calculus Lemma can lead to Logarithmic Derivative Lemma (see Theorem 5.6): for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+Let be a nonconstant meromorphic function on The Calculus Lemma can derive the Logarithmic Derivative Lemma: for any there exists a subset of finite Lebesgue measure such that
+holds for all outside (see Theorem 5.6).
 1.2. Main results
-Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper.
-Assume that is non-parabolic, by which we mean that carries a positive global Green function, i.e., it satisfies the volume growth condition
+Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and (see Section 3.2).
 Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
-The main result is the following Second Main Theorem.
+The main result is the Second Main Theorem as follows.
 Theorem 1.1 (=Theorem 6.1).
-Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
 Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
-If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
+Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 For a divisor the simple defect of with respect to is defined by
-Again, put
-Theorem 1.1 gives a defect relation:
+Set
+Since then we obtain a defect relation:
 Corollary 1.2 (=Corollary 6.3).
@@
 Let denote the scalar curvature of Set
-We obtain a Picard’s theorem:
+In the non-parabolic case, we answer the Picard’s problem as follows.
 Corollary 1.3 (=Corollary 6.10).
-Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
-Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values.
-2. Some Facts from Differential Geometry and Geometric Analysis
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
+Every meromorphic function on reduces to a constant if it omits distinct values, where denotes the maximal integer not greater than
+It implies that every meromorphic function with on reduces to a constant if it omits 2 distinct values.
+2. Some Facts from Geometric Analysis
 2.1. Volume Comparison Theorem
-A space form is defined as a complete Riemannian manifold with constant sectional curvature.
-Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Let denote the Riemannian volume of a geodesic ball centered at with radius
-The classical volume comparison theorem by Bishop-Gromov (see [5]) can be stated as follows.
+A space form is a complete Riemannian manifold with constant sectional curvature.
+Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
+The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
 Theorem 2.1 (Volume Comparison Theorem).
@@
 Corollary 2.2.
-Assume that is non-compact. If then we have
-holds for all where is the volume of a unit ball in
-When Calabi-Yau (see [35]) obtained a lower bound of
+We have
+holds for all where is the volume of the unit ball in
+When Calabi-Yau (see [39]) gave a lower bound of
 Theorem 2.3.
@@
 2.2. Estimates of Heat Kernels
-Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of is the minimal positive fundamental solution of the following heat equation
-Fix any reference point Let be the Riemannian distance function of from
+Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
 Theorem 2.4.
 Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
-holds for all
+holds for all and
 Set
-If the right-hand integral converges , then defines a minimal positive global Green function of for with a pole at i.e.,
-where is the Dirac’s delta function with a pole at It is evident that such a Green function is also unique if (1) is satisfied.
+This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
+where is the Dirac function with a pole at
+When is non-parabolic with non-negative Ricci curvature, Li-Yau [26] (see [39] also) obtained an estimate of as follows.
 Theorem 2.5.
-Assume that has non-negative Ricci curvature. If exists, then there exist constants depending only on such that
+Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
 holds for all
 3. Nevanlinna’s Functions and First Main Theorem
-Let be a non-compact complete Kähler manifold with non-negative Ricci curvature of complex dimension whose Kähler form is defined by
-in a local holomorphic coordinate Fix a reference point Let stand for the geodesic ball centered at with radius in Denote by the Riemannian volume of
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
 3.1. Construction of
-Assume that is non-parabolic, which means that satisfies the volume growth condition
-Hence, there exists a unique minimal positive global Green function of for i.e.,
-where denotes the Laplace-Beltrami operator and is the Dirac function with a pole at Let be the Riemannian distance function of from According to Li-Yau [22], there exist constants such that
-Define
-By as it is evident to see that is relatively compact for any and the sequence exhausts for According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
-which is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
-Moreover, we denote by the harmonic measure on with respect to which is defined by
+Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
+Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
+where is the heat kernel of Let be the Riemannian distance function of from Using Theorem 2.5, there exist constants such that
+holds for all For define
+By
+we see that is a bounded domain containing such that
+Furthermore, the family exhausts i.e., for any sequence with we have
+Whence, the boundary of can be formulated as
+Due to Sard’s theorem, is a submanifold of for almost all
+Set
+Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
+Let denote the harmonic measure on with respect to defined by
 where is the inward normal derivative on is the Riemannian area element of
@@
 where is the Riemannian volume element of
-Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential Since is Kählerian, then we have
-Locally, write where is a local holomorphic frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [7]), it leads to
+Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to the metrics We have
+Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
 in the sense of currents. So, we obtain the alternative expressions of and as follows
@@
 3.3. First Main Theorem
-To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) as follows.
 Lemma 3.1 (Jensen-Dynkin formula).
@@
 4. Gradient Estimates of Green Functions
-Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
-Theorem 4.1.
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
+Lemma 4.1.
 We have
@@
 then we obtain
-Let denote the gradient operator on We obtain an estimate of upper bounds of as follows:
+Let denote the gradient operator on Then, we have an upper bound of as follows.
 Theorem 4.2.
@@
 It yields from (1) that
-On the other hand, we have
-Whence, it concludes that
+Since
+then the L’Hospital’rule yields that
 where is the cut locus of and is the inward normal derivative on By
@@
 holds for all where is given by
-Proof.
 5. Calculus Lemma and Logarithmic Derivative Lemma
-Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
 5.1. Calculus Lemma
-We need the following Borel’s lemma (see, e.g., [24]):
-Lemma 5.1 (Borel Lemma).
+Lemma 5.1 (Borel’s Lemma).
 Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
@@
 Proof.
-The conclusion is true clearly for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
+The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
 exists, here is allowed. If then Set
@@
 This completes the proof. ∎
-Now, we give the Calculus Lemma:
+We establish the Calculus Lemma as follows.
 Theorem 5.2 (Calculus Lemma).
@@
 Proof.
-Invoking Theorem 4.1, we obtain
+Invoking Lemma 4.1, we have
 Set
@@
 holds for all where is given by (1). Combining the above, we have
-holds for all outside Therefore, we have the theorem proved by setting
-As a result, we conclude that
+holds for all outside Therefore, we have the theorem proved by setting ∎
 Corollary 5.3.
@@
 holds for all outside
-Proof.
 5.2. Logarithmic Derivative Lemma
 Let be a meromorphic function on The norm of the gradient is defined by
-in a local holomorphic coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
+in a holomorphic local coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
 where
@@
 On one puts a singular metric
-Then, it leads to
+which leads to
 Lemma 5.4.
@@
 Define
-Next, we establish the following Logarithmic Derivative Lemma:
+Next, we establish the following Logarithmic Derivative Lemma.
 Theorem 5.6 (Logarithmic Derivative Lemma).
@@
 Note that
-Using Lemma 5.5 and disturbing we have for any there exists a subset of finite Lebesgue measure such that
+By Lemma 5.5, we have for any there exists a subset of finite Lebesgue measure such that
 holds for all outside ∎
+By we have:
+Corollary 5.7.
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside
 6. Second Main Theorem and Defect Relation
-Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
-Set
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type.
+Recall that
 where is the Ricci form of
-Let be a divisor, where are prime divisors. The reduced form of is defined as
+Let be a divisor, where are prime divisors. The reduced form of is defined by
 6.1. Second Main Theorem
-Let be a reduced divisor of simple normal crossing type.
-We write which is an irreducible decomposition of Equip every holomorphic line bundle with a Hermitian metric which induces a Hermitian metric on Since then we can assume that As a result, gives a volume form on Pick such that and Moreover, we define a singular volume form
-on Let
-be a differentiably non-degenerate meromorphic mapping. Set
+Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
+where
+Let be a differentiably non-degenerate meromorphic mapping. Set
 It is clear that
 In further, we have
-in the sense of currents, where is the reduced form of Thus, it yields that
+in the sense of currents. Thus, it yields that
 We establish a Second Main Theorem as follows.
 Theorem 6.1 (Second Main Theorem).
-Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
 Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
-If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
+Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 Proof.
-The argument of proof is standard (see [14]), but for the completeness of this paper, we still give a proof.
-Since has only simple normal crossings, then there exist a finite open covering of and finitely many rational functions on such that are holomorphic on satisfying that
-In addition, we can require that for On write
+Since is of simple normal crossing type, then we can provide a finite open covering of and finitely many rational functions on for every such that are holomorphic on with that
+Moreover, we may require for On write
 where is a positive smooth function on Let be a partition of the unity subordinate to Set
 Again, put On we have
-Fix a one can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
-Put
+Fix any we can choose a holomorphic local coordinate near and a holomorphic local coordinate near such that
+Set
 Then, we have and at :
 Define a non-negative function on by
-Again, set for Then, at :
-That is, at :
-Combining the above, we are led to
-on Note that is bounded on whence it yields from that
-By Jensen-Dynkin formula
-Combining (6) with (7) and using Theorem 5.6, we obtain
+Moreover, put for Then, at :
+Whence, at :
+Put together the above, we are led to
+on Since is bounded on and
+then we obtain
+By Lemma 3.1
+Combining (5) with (6) and using Theorem 5.6, we have
 Using Theorem 5.2 and (4), for any there exists a subset of finite Lebesgue measure such that
-holds for all outside Thus, we conclude that
-for all outside Combining this with (6.1), we prove the theorem. ∎
-Let be a compact Riemann surface of genus Let be the holomorphic line bundle over defined by a point Note that is independent of the choice of It yields from that
+holds for all outside Hence, we conclude that
+for all outside By this with (6.1), we prove the theorem. ∎
 Corollary 6.2.
-Let be distinct points in Let be a nonconstant meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
-Let be the hyperplane line bundle over and whose Chern form is given by It yields from that
-Corollary 6.3.
-Let be hyperplanes in general position in Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
+Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
+Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 6.2. Defect Relation
-For any divisor the defect and simple defect of with respect to are defined respectively by
-Using the First Main Theorem, we see that
-For two holomorphic line bundles over we define (see [7])
+The simple defect of with respect to is defined by
+By the First Main Theorem, we have
+By the use of Theorem it is immediate that
 Corollary 6.3 (Defect Relation).
 Assume the same conditions as in Theorem Then
-Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Let denote the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
-where
-Using the definition of scalar curvature, we have
-In further, it yields that
+Let
+denote the Ricci curvature tensor of Since is Kählerian, then
+Again, denote by the scalar curvature of Note that is the trace of thus we have
+It yields that
 Corollary 6.4.
-Let be a meromorphic mapping. If there exists a small number such that satisfies the growth-curvature condition
+Let be a meromorphic mapping. If there exists a number such that satisfies the growth condition
 for sufficiently large then must be differentiably degenerate.
 Proof.
-Assume on the contrary that is differentiably non-degenerate. Then, we have as The growth-curvature condition leads to
-Thus, we obtain due to Corollary 6.4, but which contradicts with ∎
-Treat a compact Riemann surface of genus we have:
+Assume on the contrary that is differentiably non-degenerate. Then, we obtain as The growth condition implies that
+Thus, we deduce due to Corollary 6.3, but which contradicts with ∎
+Consider a compact Riemann surface of genus Note that
+for an arbitrary point where is the holomorphic line bundle over defined by Combining this with Theorem 6.1, we conclude that
 Corollary 6.5.
 Let be a nonconstant meromorphic mapping. Let be distinct points in Then
-For a meromorphic mapping
-we have
-with respect to metrics
+Let
+be a meromorphic mapping. Then
+Let be the hyperplane line bundle over Since
+then Theorem 6.1 yields that
 Corollary 6.6.
@@
 A meromorphic function on can be seen as a meromorphic mapping Set
-We obtain a Picard’s theorem:
+We obtain a Picard’s theorem as follows.
 Corollary 6.7 (Picard’s Theorem).
 Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
-6.3. Picard’s Theorem
-In the following, we investigate a more exact form of the Picard’s theorem given in Corollary 6.8. On we introduce the spherical distance
+In the following, we shall derive a better Picard’s theorem compared with Corollary 6.7. On we introduce the spherical distance
 Let be a meromorphic function on which is regarded as a meromorphic mapping
 In terms of the spherical distance, the proximity function of with respect to can be defined as
-We use the Ahlfors-Shimizu’s characteristic function of :
-Moreover, we define
-It is not hard to show the First Main Theorem:
+Adopt the Ahlfors-Shimizu’s characteristic function of as follows
+Define
+Then, we have the First Main Theorem
 Theorem 6.8 (Second Main Theorem).
-Let be a non-compact complete Kähler manifold with non-negative Ricci curvature.
-Let be distinct values in Let be a nonconstant meromorphic function on If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
+Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 Proof.
-By
-in the sense of distributions, we deduce that
-Combining this with Corollary 6.3, it gives immediately
-We re-define the simple defect of with respect to by
+Define
 Set
+Clearly, we see that
 Corollary 6.9 (Defect Relation).
 Assume the same conditions as in Theorem Then
-Consequently, we derive a more exact Picard’s theorem:
+We derive a better Picard’s theorem as follows.
 Corollary 6.10 (Picard’s Theorem).
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
 Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than

### v7

@@
-On Picard’s Theorem Via Nevanlinna Theory
+On Picard’s Problem via Nevanlinna Theory
 Abstract.
-We study the classical Picard’s problem for complete Kähler manifolds with non-negative Ricci curvature.
-Based on the global Green function method, we generalize the Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature.
-As a consequence, we give a positive answer to the Picard’s problem in the non-parabolic case, i.e., if the (non-compact) manifold is non-parabolic, then we show that every meromorphic function is a constant if it omits three distinct values.
+We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature.
+Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory.
+That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition.
 Key words and phrases:
-Nevanlinna theory; meromorphic mappings; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
-32H30; 32H25; 32A22Contents
-1. Introduction
-1.1. Motivations and Techniques
-The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces will reduce to a constant, if it omits three distinct values.
-Historically, people have been committed to developing this theorem for a long time (see [4, 16, 18, 20, 21, 24, 31]).
-The study of Picard’s problem on manifolds may have originated from the celebrated work due to S.
+Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
+32H30; 32H25; 32A221. Introduction
+1.1. Motivation
+The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values.
+Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]).
+The study of Picard’s problem on a manifold may have originated from the celebrated work due to S.
 T.
-Yau [48] on Liouville’s theorem, which proves the Liouville property of harmonic functions on non-negatively Ricci curved manifolds.
-Yau’s work has caused a classical problem: does Picard’s theorem hold on complete Kähler manifolds with non-negative Ricci curvature?
-It is well-known that Nevanlinna theory [30] is a great development of both Liouville’s theorem and Picard’s theorem.
-So, a further problem is, can Nevanlinna theory extend to complete Kähler manifolds with non-negative Ricci curvature?
-As far as we know, neither of them has been solved.
-Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see such examples in Sha-Yang [38] and Tian-Yau [43, 44]). The classical Picard’s problem states that
+Yau [46] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds.
+Yau’s work inspired people to think about the following Picard’s problem.
+Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see examples in Sha-Yang [35] and Tian-Yau [41, 42]).
+The classical Picard’s problem asks that
 Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
-Remark.
-Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem of Yau, we may have an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by By the mapping lifting theorem, admits a lifting if and only if
+Remark 1.
+Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem by Yau, one may give an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by Using the mapping lifting theorem, admits a lifting if and only if
 for some such that Since is simply-connected, then the fundamental group is trivial.
@@
 However, is not trivial in most cases.
-Thus, we cannot lift to in general.
-An early result in this direction could be dated back to 1970, S. Kobayashi [24] obtained some Picard-type theorems for holomorphic mappings between complex manifolds. In particular, he showed that
+So, we cannot lift to in general.
+An early result in this direction could be dated back to 1970, S.
+Kobayashi [21] obtained some Picard-type theorems for holomorphic mappings between certain complex manifolds.
+In particular, he showed that
 Theorem A (Kobayashi, 1970).
@@
 Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group.
-In 1975, Goldberg-Ishihara-Petridis [21] established some Picard-type theorems on the locally flat manifolds (see N.
+In 1975, Goldberg-Ishihara-Petridis [18] treated a class of locally flat manifolds (see N.
 Petridis [28] also).
@@
 Theorem B (Goldberg-Ishihara-Petridis, 1975).
-If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant mapping.
-In 2010, A.
-Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by developing a Nevanlinna-type theory based on heat diffusion, Let us introduce his main result.
-Let be the Kähler form and Laplace-Beltrami operator of associated to respectively.
-One can endow with the Fubini-Study metric Let be a meromorphic function.
-The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
-He showed that
-Theorem C (Atsuji, 2010).
-Let be a meromorphic function. Assume that satisfies the growth condition
-for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
-Atsuji’s method [2] (see [12] also) is using stochastic calculus by Brownian motion, who introduced the so-called Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [22]), he established a heat diffusion version of the Second Main Theorem which leads to the above Theorem C.
-Note that is stochastically complete (i.e., is conservative), since has non-negative Ricci curvature due to Grigor’yan’s criterion [17].
-However, in order to make and meaningful in the sense of Nevanlinna theory, the following conditions are necessary:
-for ;
-as ;
-if omits
-To do so, has to satisfy certain growth assumptions.
-That is why a growth condition is needed in Atsuji’s theorem.
-Unfortunately, the growth condition seems so strong that there are only a few meromorphic functions satisfy it.
-The main purpose of the present paper is to generalize Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature.
-It is known that Nevanlinna theory in several complex variables (see [28, 33]) studies the value distribution of meromorphic mappings between complex spaces, which has been developed for a long time.
-During this period, fruitful results have been achieved.
-For example, the reader may refer to L.
+If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant.
+It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds.
+Refer to L.
 Ahlfors [5], H.
@@
 Vojta [38] and H.
-Wu [47], etc.; and also refer to [1, 2, 3, 12, 13, 14, 15].
-To our knowledge, almost all results in Nevanlinna theory are built on complex manifolds with a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, since it is difficult to estimate the local Green function for a general bounded domain, even for a geodesic ball, under other curvature conditions.
-Recently, the author [15] gave the first result on the Second Main Theorem in Nevanlinna theory on complete Kähler manifolds with non-negative Ricci curvature by introducing the global Green function method.
-Consequently, the author answered the Picard’s problem affirmatively provided that is of maximal volume growth:
-where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
-The original technique in [15] is to construct a family of relatively compact exhaustive domains of using the minimal positive global Green function for Based on an asymptotic estimate of obtained by Colding-Minicozzi [9], one can obtain two-sided bounds of the local Green function for in terms of integral forms.
-As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce his main result as follows.
-Colding-Minicozzi [9] (see [25] also) described the asymptotic behavior of : there exists a constant such that
-where denotes the Riemannian distance function of from By means of the asymptotic estimate, the author defined a family of relatively compact domains as follows
-for Hence, the Green function of for with a pole at satisfying Dirichlet boundary condition can be formulated as
-According to Lemma 3.5 in [15], for any there exists such that
-holds for all satisfying Based on this estimate, the Calculus Lemma was established.
-For is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [25]).
-So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [15], Theorem 3.10), which gives an immediate result:
-Theorem D (Dong, 2023).
-If is of maximal volume growth, then every meromorphic function on must be a constant if it omits distinct values.
-However, the technique in [15] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no such asymptotic estimates for if is not of maximal volume growth.
-In this paper, we will generalize the Nevanlinna theory in the case that is a non-parabolic manifold, i.e., there exists a positive global Green function for Note that is non-parabolic if is of maximal volume growth.
-By N.
-Varopoulos [45, 46] (see Li-Tam-Wang [23, 25] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted also by Our original idea is the construction of domain and application of Li-Yau’s estimate on heat kernel and global Green function.
-Using Li-Yau’s estimate [26], there exist constants such that
-holds for all and For we redefine
-Then, we have
-which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and : for we have
-which plays an important role in the establishment of Nevanlinna theory in this paper (see Lemma 4.1).
-Negative Curvature Method and Logarithmic Derivative Lemma Method are very typical approaches in the study of Nevanlinna theory.
-In this paper, the Logarithmic Derivative Lemma Method is used.
-So, we need to establish the so-called Calculus Lemma which is a way to the Logarithmic Derivative Lemma.
-We will introduce two techniques in the paper: Gradient Estimates (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
-Gradient Estimates. For any there exists such that
-holds for all
-The above gradient estimate yields an estimate of the harmonic measure on with respect to : for any there exists such that
-holds for all (see Corollary 4.3).
-Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
-holds for all outside
-Let be a nonconstant meromorphic function on The Calculus Lemma can derive the Logarithmic Derivative Lemma: for any there exists a subset of finite Lebesgue measure such that
-holds for all outside (see Theorem 5.6).
+Wu [45], and see [1, 2, 3, 10, 11, 12] and references therein.
+In 2010, A.
+Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by establishing a Nevanlinna-type theory based on heat diffusion, We shall introduce his main work.
+Let be the Kähler form and Laplace-Beltrami operator of associated to respectively.
+One can endow with the Fubini-Study metric Let be a meromorphic function.
+The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
+He showed that
+Theorem C (Atsuji, 2010).
+Let be a meromorphic function. Assume that satisfies the growth condition
+for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
+Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion.
+Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and utilizing Itô’s formula (see [19]), he established an analogue of Nevanlinna theory.
+Note that is stochastically complete (i.e., is conservative) since Grigor’yan’s criterion [14].
+To make and meaningful, the following conditions are necessary:
+for ;
+as ;
+if omits
+To do so, has to satisfy certain growth assumption. That is why a growth condition is needed in Theorem C.
+In this paper, we generalize Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature.
+Our innovative method is to construct a family of exhaustive precompact domains for by means of a unique minimal positive global Green function for so that we can well define the Nevanlinna’s functions on and establish a Second Main Theorem by introducing the global Green function method.
+Apply the Second Main Theorem, we can give a positive answer to the Picard’s problem under certain condition.
 1.2. Main results
-Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and (see Section 3.2).
-Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
-The main result is the Second Main Theorem as follows.
+Assume that is non-parabolic, by which we mean that there exists a positive global Green function for By N.
+Varopoulos [43, 44] (see Li-Tam-Wang [20, 22] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted by where is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
+holds for all and where is the Riemannian distance function of from and denotes the Riemannian volume of the geodesic ball centered at with radius in For define
+Then, for all and the family exhausts
+Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and on (see Section 3.2).
+Let denote the Ricci form of associated to the metric and be the canonical line bundle over The characteristic function of is defined by
+Again, set
+The main result is the following Second Main Theorem.
 Theorem 1.1 (=Theorem 6.1).
@@
 Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+holds for all outside where is given by
 For a divisor the simple defect of with respect to is defined by
 Set
-Since then we obtain a defect relation:
+We obtain a defect relation:
 Corollary 1.2 (=Corollary 6.2).
 Assume the same conditions as in Theorem Then
-Let denote the scalar curvature of Set
-In the non-parabolic case, we answer the Picard’s problem as follows.
-Corollary 1.3 (=Corollary 6.10).
+if one of the following conditions is satisfied
+satisfies the volume growth condition
+is of non-polynomial type growth, i.e., satisfies the growth condition
+Remark 2. Note that due to
+The condition (2) is relaxed. For example, it holds when with and holds when with or and
+As an application of Corollary 1.2, we answer the Picard’s problem under certain condition as follows.
+Corollary 1.3.
 Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
-Every meromorphic function on reduces to a constant if it omits distinct values, where denotes the maximal integer not greater than
-It implies that every meromorphic function with on reduces to a constant if it omits 2 distinct values.
+If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
+Corollary 1.4.
+Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
 2. Some Facts from Geometric Analysis
@@
 The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
-Theorem 2.1 (Volume Comparison Theorem).
+Theorem 2.1.
 If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
 holds for all
-In particular, if then we obtain:
+When we obtain:
 Corollary 2.2.
@@
 holds for all where is the volume of the unit ball in
-When Calabi-Yau (see [39]) gave a lower bound of
+When Calabi-Yau (see [36]) proved that
 Theorem 2.3.
@@
 holds for all
-2.2. Estimates of Heat Kernels
+2.2. Estimates of Heat Kernel
 Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
@@
 This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
-where is the Dirac function with a pole at
-When is non-parabolic with non-negative Ricci curvature, Li-Yau [26] (see [39] also) obtained an estimate of as follows.
+where is the Dirac’s delta function with a pole at
+When is non-parabolic with non-negative Ricci curvature, Li-Yau [23] (see [36] also) gave two-sided bounds of as follows.
 Theorem 2.5.
@@
 By
-we see that is a bounded domain containing such that
-Furthermore, the family exhausts i.e., for any sequence with we have
+where means we derive that is a precompact domain containing such that
+In fact, the family exhausts namely, for any sequence with we have
 Whence, the boundary of can be formulated as
@@
 where is the Riemannian volume element of
-Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to the metrics We have
 Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
@@
 Since
-then we obtain
-Let denote the gradient operator on Then, we have an upper bound of as follows.
+we obtain
+Let denote the gradient operator on a Riemannian manifold. Cheng-Yau [9] proved the following theorem.
+Lemma 4.2.
+Let be a complete Riemannian manifold. Let be a non-negative harmonic function on a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
+where is the Riemannian distance between and and is the lower bound of the Ricci curvature on
+We have an upper bound of as follows.
 Theorem 4.3.
-For any there exists such that
-holds for all where is given by
+There exists a constant such that
 Proof.
-It yields from (1) that
-Since
-then the L’Hospital’rule yields that
-where is the cut locus of and is the inward normal derivative on By
-we see that for any there exists such that
-holds for all Since has measure 0, with the aid of the continuity of on and on one then has the theorem proved. ∎
-As a result, we obtain an estimate for upper bounds of harmonic measure on with respect to as follows:
+By
+we can obtain an estimate for upper bounds of as follows.
 Corollary 4.4.
-For any there exists such that
-holds for all where is given by
+There exists a constant such that
+where is the Riemannian area element of
 5. Calculus Lemma and Logarithmic Derivative Lemma
@@
 exists, here is allowed. If then Set
-Note that exists for almost every Then, we have
+Note that exists for almost all Then, we have
 This completes the proof. ∎
@@
 Theorem 5.2 (Calculus Lemma).
-Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
-holds for all outside
+Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is given by
 Proof.
@@
 In further, we have
-Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
-holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
-holds for all where is given by (1). Combining the above, we have
-holds for all outside Therefore, we have the theorem proved by setting ∎
-Corollary 5.3.
-Let be a locally integrable function on Assume that is locally bounded at Assume that is of maximal volume growth. Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant such that
+Set Combining the above, we have
+holds for all outside ∎
 5.2. Logarithmic Derivative Lemma
 Let be a meromorphic function on The norm of the gradient is defined by
-in a holomorphic local coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
+in holomorphic local coordinates here is the inverse of Again, define the Nevanlinna’s characteristic function of by
 where
@@
 Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+holds for all outside where is given by
 Proof.
 Define
-Next, we establish the following Logarithmic Derivative Lemma.
+We establish the following Logarithmic Derivative Lemma.
 Theorem 5.5 (Logarithmic Derivative Lemma).
 Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+holds for all outside where is given by
 Proof.
@@
 By Lemma 5.4, we have for any there exists a subset of finite Lebesgue measure such that
-holds for all outside ∎
-By we have:
-Corollary 5.7.
-Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+holds for all outside where is given by ∎
 6. Second Main Theorem and Defect Relation
@@
 Let be a divisor, where are prime divisors. The reduced form of is defined by
-6.1. Second Main Theorem
 Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
@@
 Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
+holds for all outside where is given by
 Proof.
-Since is of simple normal crossing type, then we can provide a finite open covering of and finitely many rational functions on for every such that are holomorphic on with that
-Moreover, we may require for On write
+Since is of simple normal crossing type, then there are a finite open covering of and finitely many rational functions on for each such that are holomorphic on with that
+In addition, we can require that for all On write
 where is a positive smooth function on Let be a partition of the unity subordinate to Set
 Again, put On we have
-Fix any we can choose a holomorphic local coordinate near and a holomorphic local coordinate near such that
+Fix any one can take holomorphic local coordinates near and holomorphic local coordinates near such that
 Set
@@
 holds for all outside Hence, we conclude that
-for all outside By this with (6.1), we prove the theorem. ∎
-Corollary 6.2.
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
-Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
-Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
-6.2. Defect Relation
-The simple defect of with respect to is defined by
+for all outside By this with (6), we prove the theorem. ∎
+Recall that the simple defect of with respect to is defined by
 By the First Main Theorem, we have
@@
 Assume the same conditions as in Theorem Then
-Let
-denote the Ricci curvature tensor of Since is Kählerian, then
-Again, denote by the scalar curvature of Note that is the trace of thus we have
-It yields that
-Corollary 6.4.
-Let be a meromorphic mapping. If there exists a number such that satisfies the growth condition
-for sufficiently large then must be differentiably degenerate.
+if one of the following conditions is satisfied
+satisfies the volume growth condition
+is of non-polynomial type growth, i.e., satisfies the growth condition
 Proof.
-Assume on the contrary that is differentiably non-degenerate. Then, we obtain as The growth condition implies that
-Thus, we deduce due to Corollary 6.3, but which contradicts with ∎
-Consider a compact Riemann surface of genus Note that
-for an arbitrary point where is the holomorphic line bundle over defined by Combining this with Theorem 6.1, we conclude that
-Corollary 6.5.
-Let be a nonconstant meromorphic mapping. Let be distinct points in Then
-Let
-be a meromorphic mapping. Then
-Let be the hyperplane line bundle over Since
-then Theorem 6.1 yields that
-Corollary 6.6.
-Let be a differentiably non-degenerate meromorphic mapping. Let be hyperplanes in general position in Then
-A meromorphic function on can be seen as a meromorphic mapping Set
-We obtain a Picard’s theorem as follows.
-Corollary 6.7 (Picard’s Theorem).
-Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
-In the following, we shall derive a better Picard’s theorem compared with Corollary 6.7. On we introduce the spherical distance
-Let be a meromorphic function on which is regarded as a meromorphic mapping
-In terms of the spherical distance, the proximity function of with respect to can be defined as
-Adopt the Ahlfors-Shimizu’s characteristic function of as follows
-Define
-Then, we have the First Main Theorem
-Theorem 6.8 (Second Main Theorem).
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
-Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for all outside
-Proof.
-Define
-Set
-Clearly, we see that
-Corollary 6.9 (Defect Relation).
-Assume the same conditions as in Theorem Then
-We derive a better Picard’s theorem as follows.
-Corollary 6.10 (Picard’s Theorem).
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
-Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
+Note that due to Corollary 2.2. We have
+Using the volume growth condition, we obtain
+Since a non-constant meromorphic mapping has the growth rate at least, then we have the theorem proved if is satisfied due to Theorem 6.1.
+For we have
+where
+So, it yields from that
+This completes the proof. ∎

### v8

@@
 Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory.
-That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition.
+That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition; and prove that every meromorphic function of non-polynomial type growth on such a manifold can avoid 2 distinct values at most.
 Key words and phrases:
-Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
+Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2020 Mathematics Subject Classification:
 32H30; 32H25; 32A221. Introduction
@@
 Yau [47] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds.
-Yau’s work inspired people to think about the following Picard’s problem.
-Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see examples in Sha-Yang [35] and Tian-Yau [41, 42]).
-The classical Picard’s problem asks that
+Yau’s work inspires people to think about the following Picard’s problem.
+Let be a complete noncompact Kähler manifold with non-negative Ricci curvature (see examples for such manifolds in Sha-Yang [35] and Tian-Yau [40, 41]).
+The classical Picard’s problem states that
 Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
@@
 Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem by Yau, one may give an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by Using the mapping lifting theorem, admits a lifting if and only if
-for some such that Since is simply-connected, then the fundamental group is trivial.
-It implies that is necessarily a trivial group if such a lifting exists.
+for some such that Since is simply-connected, the fundamental group is trivial.
+This implies that is necessarily a trivial group if such a lifting admits.
 However, is not trivial in most cases.
-So, we cannot lift to in general.
+Thus, we cannot lift to in general.
 An early result in this direction could be dated back to 1970, S.
@@
 Theorem A (Kobayashi, 1970).
-If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
+Assume that there is a complex Lie group acting on transitively. Then, every meromorphic function on must be a constant if it omits distinct values.
 Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group.
@@
 It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds.
-Refer to L.
+For instance, refer to L.
 Ahlfors [5], H.
@@
 Stoll [38, 39], P.
-Vojta [38] and H.
-Wu [45], and see [1, 2, 3, 10, 11, 12] and references therein.
+Vojta [45], H.
+Wu [46], and also refer to [1, 2, 3, 10, 11, 12] and references therein.
 In 2010, A.
-Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by establishing a Nevanlinna-type theory based on heat diffusion, We shall introduce his main work.
-Let be the Kähler form and Laplace-Beltrami operator of associated to respectively.
-One can endow with the Fubini-Study metric Let be a meromorphic function.
-The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
+Atsuji [2] gave an affirmative answer for a class of meromorphic functions with slower growth by introducing a Nevanlinna-type theory based on heat diffusion, We introduce his main work.
+Let stand for the Kähler form and Laplace-Beltrami operator on respectively.
+Let be a meromorphic function.
+We equip with Fubini-Study metric The Kählerness of indicates that the Hilbert-Schmidt norm of differential with respect to metrics can be expressed as
 He showed that
@@
 Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion.
-Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and utilizing Itô’s formula (see [19]), he established an analogue of Nevanlinna theory.
-Note that is stochastically complete (i.e., is conservative) since Grigor’yan’s criterion [14].
-To make and meaningful, the following conditions are necessary:
+Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [19]), he established an analogue of the Nevanlinna theory.
+Note that is stochastically complete (i.e., is conservative) due to Grigor’yan’s criterion (see [14]).
+In order to make and meaningful, the following are necessary:
 for ;
@@
 if omits
-To do so, has to satisfy certain growth assumption. That is why a growth condition is needed in Theorem C.
-In this paper, we generalize Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature.
-Our innovative method is to construct a family of exhaustive precompact domains for by means of a unique minimal positive global Green function for so that we can well define the Nevanlinna’s functions on and establish a Second Main Theorem by introducing the global Green function method.
-Apply the Second Main Theorem, we can give a positive answer to the Picard’s problem under certain condition.
+Hence, needs to satisfy certain growth assumption. That is why a growth condition is added to Theorem C.
+In this paper, our purpose is to study the Picard’s problem using the tool of Nevanlinna theory, i.e., we shall extend the classical Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature, and apply the Second Main Theorem to establish a Picard’s theorem.
+Let us first recall the basic notions of the non-parabolicity and the Ricci curvature of Riemannian manifolds.
+Let be a complete Riemannian manifold.
+One says that is non-parabolic, if there exists a positive global Green function for and parabolic otherwise.
+Or equivalently, in a viewpoint of geometric analysis, is called non-parabolic if
+for and parabolic if this infinite integral is divergent, where denotes the transition density function of the Brownian motion generated by the Laplace-Beltrami operator on and which is also called the heat kernel of Note that when is non-parabolic, defines the unique minimal positive global Green function of for More details may refer to [20, 22, 42, 43].
+Let stand for the Riemannian curvature tensor of The Ricci curvature tensor of is defined by
+for any with where is an orthonormal basis of The Ricci curvature of at in a direction is defined by
+We say that is of non-negative Ricci curvature at if for each nonzero vector and that is of non-negative Ricci curvature, if it is of non-negative Ricci curvature at each Moreover, the sectional curvature of at along a section spanned by is defend by
+From the definitions, the sign of the sectional curvature at a point determines the sign of the Ricci curvature at the same point.
+To facilitate the construction of examples later for manifolds considered in this paper, we shall provide some criteria of non-parabolicity via the volume growth.
+A sharp necessary condition by N.
+Varopoulos [44] states that if is non-parabolic, then
+where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in However, (1) is far from sufficient, and a counterexample was constructed in [44].
+The first major result for the sufficiency was due to N.
+Varopoulos [43] and Li-Yau [23].
+Based on the heat kernel estimates, they proved that if has non-negative Ricci curvature and (1) is satisfied, then is non-parabolic.
+So, for with non-negative Ricci curvature, is non-parabolic if and only if (1) is satisfied.
+Next, assume that is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Here, the non-parabolicity of is meant that is non-parabolic as a Riemannian manifold.
+The Chern-Ricci form of on is defined by
+with Since is Kähler, is equivalent to Our innovative idea in the present paper is to construct a family of exhaustive precompact domains for by means of the minimal positive global Green function for so that it allows us to well define Nevanlinna’s functions on and then establish a Second Main Theorem by estimating the gradient of Green function for Applying the Second Main Theorem, we provide a positive answer to the Picard’s problem under certain condition.
 1.2. Main results
-Assume that is non-parabolic, by which we mean that there exists a positive global Green function for By N.
-Varopoulos [43, 44] (see Li-Tam-Wang [20, 22] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted by where is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
-holds for all and where is the Riemannian distance function of from and denotes the Riemannian volume of the geodesic ball centered at with radius in For define
-Then, for all and the family exhausts
-Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and on (see Section 3.2).
-Let denote the Ricci form of associated to the metric and be the canonical line bundle over The characteristic function of is defined by
-Again, set
+Let be the minimal positive global Green function of for in which is the Laplace-Beltrami operator on and is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
+holds for all and all in which is the Riemannian distance between and and stands for the Riemannian volume of the geodesic ball centered at with radius in For define
+Then, we have for all and the family exhausts
+Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions on (see Section 3.2).
+Let be the canonical line bundle over The characteristic function of is defined by
+where is the Kähler form of and is the Green function of for with a pole at satisfying Dirichlet boundary condition. Set
 The main result is the following Second Main Theorem.
@@
 is of non-polynomial type growth, i.e., satisfies the growth condition
-Remark 2. Note that due to
+Remark 2. Note that as due to
 The condition (3) is relaxed. For example, it holds when with and holds when with or and
-As an application of Corollary 1.2, we answer the Picard’s problem under certain condition as follows.
+As an application of Corollary 1.2, we consider the Picard’s problem under certain condition below.
+Treat the case when with the point line bundle, i.e., the hyperplane line bundle Take where are distinct points in If we put then A basic and well-known fact states that
+which yields that
+whenever or in Corollary 1.2 is satisfied. According to the arguments above, we conclude that
 Corollary 1.3.
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature.
 If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
 Corollary 1.4.
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
+To close the section, we finally provide some typical examples for where Corollaries 1.3 and 1.4 apply.
+According to (1), we first notice that with is a non-parabolic complete noncompact Kähler manifold with zero Ricci curvature.
+Below, we give two nontrivial examples.
+Example 1.
+For an integer we equip with Fubini-Study metric Note that is a complete Kähler manifold with positive sectional curvature (and thus with positive Ricci curvature).
+We consider the product manifold with which is equipped with the product metric induced from and the standard Euclidean metric on Because is a complete noncompact Kähler manifold, so is Since is flat, is clearly Ricci non-negatively curved under this product metric.
+Again, by the dimension condition that we deduce that is of the volume growth: as which leads to that is satisfied.
+Therefore, is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature.
+It is not difficult to see that there exist a lot of nonconstant meromorphic functions on For instance, we can take or where is any meromorphic function on and is any rational function on
+Example 2.
+Treat a product manifold with and where denotes the -dimensional complex torus equipped with the metric inherited from the standard Euclidean metric on Note that is a complete noncompact Kähler manifold with zero Ricci curvature under the induced product metric.
+By the similar argument as in Example 1 also shows the non-parabolicity of The existence of nonconstant meromorphic functions on can be confirmed easily.
+For instance, for we can take or where is any meromorphic function on and is any elliptic function with a periodic lattice on
+More general, can be taken as the product of and in which is any complete compact Kähler manifold with non-negative Ricci curvature, and is a complete noncompact Kähler manifold with non-negative Ricci curvature and volume growth satisfying (1).
+In particular, satisfies the desired conditions.
 2. Some Facts from Geometric Analysis
 2.1. Volume Comparison Theorem
-A space form is a complete Riemannian manifold with constant sectional curvature.
-Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
+A space form is a complete (simply-connected) Riemannian manifold with constant sectional curvature.
+Let denote the -dimensional space form with constant sectional curvature and let denote the Riemannian volume of a geodesic ball with radius in Let denote a complete Riemannian manifold of dimension with Ricci curvature tensor and let denote the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
 The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
 Theorem 2.1.
-If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
+If then the volume ratio is a non-increasing function in and which tends to as Hence, we have
 holds for all
@@
 We have
-holds for all where is the volume of the unit ball in
-When Calabi-Yau (see [36]) proved that
+holds for all where denotes the standard Euclidean volume of the unit ball in
+When Calabi-Yau (see [36]) showed that
 Theorem 2.3.
-Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
+Assume that is noncompact. If then has an infinite volume. More precisely, for any there exists a constant such that
 holds for all
-2.2. Estimates of Heat Kernel
+2.2. Estimates of Heat Kernels
 Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
+Li-Yau [23] (see [36] also) obtained the two-sided estimates of
 Theorem 2.4.
 Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
-holds for all and
+holds for all and all
 Set
@@
 Theorem 2.5.
-Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
+Assume that has non-negative Ricci curvature. If is non-parabolic, then there exist constants depending only on such that
 holds for all
 3. Nevanlinna’s Functions and First Main Theorem
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to metric is defined by
 3.1. Construction of
-Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
+Fix a reference point Let be the Riemannian volume of where denotes the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
 Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
@@
 holds for all For define
-By
-where means we derive that is a precompact domain containing such that
-In fact, the family exhausts namely, for any sequence with we have
-Whence, the boundary of can be formulated as
-Due to Sard’s theorem, is a submanifold of for almost all
+Since
+one can conclude immediately that is a precompact domain containing such that whenever and that
+Thus, the family exhausts i.e., for any sequence with we have
+The boundary of can be described as
+By Sard’s theorem, is smooth for almost all
 Set
-Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
-Let denote the harmonic measure on with respect to defined by
-where is the inward normal derivative on is the Riemannian area element of
+Evidently, is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
+where is the Dirac’s delta function with a pole at Let stand for the harmonic measure on with respect to defined by
+where is the inward normal derivative on and is the induced Riemannian area element of
 3.2. Nevanlinna’s Functions
-In what follows, we will introduce Nevanlinna’s functions.
-Let be a meromorphic mapping, where is a complex projective manifold.
-Let be a Hermitian holomorphic line bundle over with the Chern form where
+In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form
+where
+so that
 Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
 where is the Riemannian volume element of
-Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
+Locally, write in which is a local holomorphic frame of and is a holomorphic function. Using Poincaré-Lelong formula (see, e.g., [8]), we get
 in the sense of currents. So, we obtain the alternative expressions of and as follows
@@
 3.3. First Main Theorem
-Lemma 3.1 (Jensen-Dynkin formula).
-Let be a -class function on outside a polar set of singularities at most. Assume that Then
-Assume that Apply Jensen-Dynkin formula to then we are led to
-Therefore, we are led to that
+Lemma 3.1 (Green-Dynkin formula).
+Let be a function on outside a polar set of singularities at most. Assume that Then
+Assume that Apply Green-Dynkin formula to we are led to
+To conclude, we obtain:
 Theorem 3.2 (First Main Theorem).
@@
 4. Gradient Estimates of Green Functions
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
+To give a gradient estimate of we need to introduce some lemmas as follows.
 Lemma 4.1.
@@
 According to the definition of Green function for it is immediate that for
-Since
-we obtain
-Let denote the gradient operator on a Riemannian manifold. Cheng-Yau [9] proved the following theorem.
+Since for we obtain
+Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [9] proved the following theorem.
 Lemma 4.2.
-Let be a complete Riemannian manifold. Let be a non-negative harmonic function on a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
-where is the Riemannian distance between and and is the lower bound of the Ricci curvature on
-We have an upper bound of as follows.
+Let be a complete Riemannian manifold of dimension Let be any geodesic ball centered at with radius in Then, there exists a constant depending only on such that
+holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
+We obtain an upper estimate of as follows.
 Theorem 4.3.
-There exists a constant such that
+There exists a constant independent of such that
 Proof.
 By
-we can obtain an estimate for upper bounds of as follows.
+we can derive an upper estimate of as follows.
 Corollary 4.4.
-There exists a constant such that
-where is the Riemannian area element of
+There exists a constant independent of such that
+where is the induced Riemannian area element of
 5. Calculus Lemma and Logarithmic Derivative Lemma
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
 5.1. Calculus Lemma
-Lemma 5.1 (Borel’s Lemma).
-Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
+Lemma 5.1 (Borel’s Growth Lemma).
+Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 Proof.
-The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
-exists, here is allowed. If then Set
+The conclusion is clearly true for Next, we assume that Since is non-decreasing, there is a number such that Using the non-decreasing property of the limit exists or If then Set
 Note that exists for almost all Then, we have
 This completes the proof. ∎
-We establish the Calculus Lemma as follows.
+We establish the following Calculus Lemma.
 Theorem 5.2 (Calculus Lemma).
-Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
+Let be a locally integrable function on Assume that is locally bounded at Then there is a constant independent of such that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside where is given by
@@
 In further, we have
-Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
-holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant such that
+Employing Borel’s growth lemma to the left hand side of the above equality twice: one is to and another is to then we conclude that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant independent of such that
 Set Combining the above, we have
-holds for all outside ∎
+holds for all outside The proof is completed. ∎
 5.2. Logarithmic Derivative Lemma
 Let be a meromorphic function on The norm of the gradient is defined by
-in holomorphic local coordinates here is the inverse of Again, define the Nevanlinna’s characteristic function of by
+in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
 where
-It is not difficult to show that
-On one puts a singular metric
-which leads to
+For any it is not hard to deduce that
+Put a singular metric on :
 Lemma 5.3.
@@
 Proof.
-Note that
-Whence, it concludes from Fubini’s theorem that
+A direct computation gives
+Whence, we conclude from Fubini’s theorem that
 Lemma 5.4.
@@
 Define
-We establish the following Logarithmic Derivative Lemma.
+With the above preparations, we are to prove the following Logarithmic Derivative Lemma.
 Theorem 5.5 (Logarithmic Derivative Lemma).
@@
 Proof.
-Note that
+The concavity of gives that
 By Lemma 5.4, we have for any there exists a subset of finite Lebesgue measure such that
@@
 6. Second Main Theorem and Defect Relation
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type.
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be a complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type.
 Recall that
-where is the Ricci form of
+where is the Chern-Ricci form of on
 Let be a divisor, where are prime divisors. The reduced form of is defined by
 Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
-where
 Let be a differentiably non-degenerate meromorphic mapping. Set
@@
 in the sense of currents. Thus, it yields that
-We establish a Second Main Theorem as follows.
+We give the main theorem in this paper, i.e., the Second Main Theorem as follows.
 Theorem 6.1 (Second Main Theorem).
-Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature.
+Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature.
 Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping.
@@
 Proof.
-Since is of simple normal crossing type, then there are a finite open covering of and finitely many rational functions on for each such that are holomorphic on with that
+Since has the simple normal crossing type, there exist a finite open covering of and finitely many rational functions on for each such that are holomorphic on with
 In addition, we can require that for all On write
@@
 Again, put On we have
-Fix any one can take holomorphic local coordinates near and holomorphic local coordinates near such that
-Set
-Then, we have and at :
+Fix any Take a local holomorphic coordinate near and a local holomorphic coordinate near such that
+and
+Set Then, we have and
 Define a non-negative function on by
-Moreover, put for Then, at :
-Whence, at :
+Moreover, put for Then
+which yields that
 Put together the above, we are led to
 on Since is bounded on and
-then we obtain
+we obtain
 By Lemma 3.1
@@
 Proof.
-Note that due to Corollary 2.2. We have
-Using the volume growth condition, we obtain
-Since a non-constant meromorphic mapping has the growth rate at least, then we have the theorem proved if is satisfied due to Theorem 6.1.
-For we have
+We first show that the corollary holds if is satisfied. By Corollary 2.2, we have Thus, we are led to that
+Using the condition we obtain
+Since each non-constant meromorphic mapping has growth at least, we have the theorem proved if is satisfied due to Theorem 6.1. For we obtain
 where
-So, it yields from that
+Thus, it yields from that
 This completes the proof. ∎